OpenStudy (anonymous):

How can I multiply the price of gasoline 3.929 by 10%? Would I multiply it by 1.10?

OpenStudy (tkhunny):

I can think of at least three things you might mean. Can you explain what it is you are doing?

OpenStudy (anonymous):

I need to multiply the price of gas by 10%, the current price is 3.929. To multiply it by 10% should I multiply the price by 1.10?

OpenStudy (tkhunny):

You didn't say anything new. Do you have an actual problem statement? Gas prices increase 10%: 3.929 * (1 + 0.10) = ?? Gas prices decrease 10%: 3.929 * (1 - 0.10) = ?? We just want to know 10% of 3.929: 3.929*0.10 = ??

OpenStudy (anonymous):

thanks!

OpenStudy (tkhunny):

Sure. Not sure what I did, but okay!